My sister is 2 years older than I am. The sum of our ages is 18. What is our ages?
Answer by stanbon(75887)

My sister is 2 years older than I am.
The sum of our ages is 18. What is our ages?
---------------------
Equations:
S = I + 2
S + I = 18
---------------
Substitute for "S" and solve for "I":
I + 2 + I = 18
----
2I = 16
I = 8 yrs (I's age)
-----
Solve for "S":
S = I + 2
S = 8 + 2 = 10 yrs (Sis's age)
-------
Cheers,
Stan H.
==============